1) A school has ordered some uniforms for students of years 7, 8, 9. Year 7 gets 42% of the uniforms; year 8 gets 5/7 of the uniforms that year 7 gets;year 9 gets 20 uniforms less than year 8. Find the total number of uniforms ordered.

Answer by macston(5194)

Year 7=42%
Year 8=(5/7)(42%)=30%
Year 9=100%-72%=28%
Year 9=Year 8-20=30%-20uniforms 30%-20 uniforms=28% so 2%=20 uniforms
Total (100%)=20/0.02=1000 uniforms
ANSWER: The total order was 1000 uniforms.